If a solution is 0.001 M in NaOH, what is the pH?

a) 3

b) 4

c) 7

d) 11
1
Expert's answer
2014-06-17T07:06:04-0400
NaOH completely dissociates in water solution:
NaOH --> Na+ + OH-
Therefore, the concentration of OH- ions is equal to the concentration of the solution:
c(OH-) = c(NaOH) = 0.001
According to water self-ionization equation,
Kw = c(H+)c(OH-) = 10-14
Hence,
c(H+) = Kw/c(OH-) = 10-14/10-3 = 10-11
pH = - lg c(H+) = 11
So, the correct option is (d): 11